integrate xcube/x+2

Grade:11

2 Answers

Roshan Mohanty
10 years ago

int(x^3 / x+2) = int ( x^3 + 8 - 8 /(x+2))
= int([x+2][x^2 - 2x +4 ] / [x+2]) - int(8/[x+2])
= int(x^2 - 2x +4) - 8ln(x+2) +c
=x^3/3 - x^2 + 4x - 8ln(x+2) + C

Plzzz approve :) 

Akash Kumar Dutta
10 years ago

int x^3/x+2
=int x^2-2x+4 - 8/x+2
=x^3/3 -x^2 +4x-8ln|x+2| +C